Jeff wants to put 2/5 of his earnings into his savings account. If he worked for 15 hours and made $9 per hour how much should he place into his savings

Answer by checkley79(3341)

9*15=$135 AMOUNT EARNED.
(2/5)*135=270/5=$54 AMOUNT TO BE SAVED.